If you have 72 mL of a 0.67 M solution and wanted to use it to make a 0.012 M solution, what would the total volume (in mL) of the new solution be?
Answer #1

M1V1=M2V2

0.67(M)*72(m)l=0.012(M)*V2(ml)

V2(ml)=(0.67(M)*72(ml))/0.012(M)

V2=2040ml.

therefore the required total volume of the new solution is 2040ml.